# Dissecting the Hubble tension: Insights from a diverse set of Sound Horizon-free measurements

### v1

@@ II.5 Category 3: Pure Local One-Step (Model Independent)
| Kuo [Kuo2013] | 2013 | 68.0 | 9.0 | Megamasers (MCP) |
| Gao [Gao2015] | 2015 | 66.0 | 6.0 | Megamasers |
| Zhang [Zhang2022] | 2022 | 65.9 | 3.0 | CC + HII |
| Du [Du2023] | 2023 | 71.5 | 3.8 | TD Lenses + GRBs |
| Sneppen [Sneppen2023] | 2023 | 67.0 | 3.6 | EPM + Kilonovae |
| Li [Li2024a] | 2024 | 66.3 | 3.7 | Lensing Bias Correction |
| Vogl [Vogl2024] | 2024 | 74.9 | 2.7 | SNe II (Tailored EPM) |
| Song [Song2025] | 2025 | 70.4 | 6.9 | GW + SGL Calibration |
| Favale [Favale2025] | 2025 | 68.8 | 3.0 | Cosmic Chronometers |
We highlight that this category exhibits significant internal heterogeneity. Notably, the Vogl [Vogl2024] tailored EPM measurement ( km s-1 Mpc-1) strongly favor the Distance Ladder value, while most cosmic chronometer measurements favor lower values. This heterogeneity will be discussed further in Section IV.
II.6 Category 4: CMB Sound Horizon Free
The final category includes 6 measurements that utilize early-Universe data (primarily CMB) but specifically avoid the Sound Horizon as a standard ruler. Instead, these methods rely on alternative physical scales that are calibrated independently of pre-recombination acoustic physics:
Matter-radiation equality scale (): The comoving wavenumber corresponding to the horizon size at matter-radiation equality provides a standard ruler that depends on and but is independent of [Philcox2022, Bahr-Kalus2025].
BAO without priors: Analyses that use BAO peak positions but treat as a free parameter, constraining through the shape of rather than absolute calibration [Pogosian2024].
Sunyaev-Zel’dovich (SZ) effect: Combining SZ decrements with X-ray observations of galaxy clusters to infer angular diameter distances independent of [Reese2003, Colaco2023].
@@ II.6 Category 4: CMB Sound Horizon Free
| Reese [Reese2003] | 2003 | 61.0 | 18.0 | SZ Effect |
| Colaco [Colaco2023] | 2023 | 67.2 | 6.1 | SZ + X-ray |
| Pogosian [Pogosian2024] | 2024 | 69.5 | 0.94 | BAO-DESI (No ) |
| Pogosian [Pogosian2024] | 2024 | 68.0 | 0.94 | BAO-pre DESI (No ) |
| Bahr-Kalus [Bahr-Kalus2025] | 2025 | 65.2 | 5.6 | DESI Turnover Scale |
| Krolewski [Krolewski2025] | 2025 | 69.0 | 2.5 | DESI + CMB (No ) |
We note that this category includes two entries from Pogosian et al. [Pogosian2024], corresponding to analyses using DESI and pre-DESI BAO data respectively. While these share a common methodology, they utilize substantially different datasets and thus provide partially independent constraints. However, the very low reduced chi-squared () for this category suggests that uncertainties may be conservative or that some correlation exists between measurements. We address this in the correlation analysis of Appendix C.
@@ II.7 Summary of Classification
| 1 | Distance Ladder | 29 | Yes | No |
| 2 | Local One-Step (CDM) | 33 | No | Yes |
| 3 | Pure Local One-Step | 15 | No | No |
| 4 | CMB Sound Horizon Free | 6 | No | Yes∗ |
@@ III Statistical Analysis and Results
In this section, we present a detailed statistical analysis of the measurements classified in Section II. We utilize the inverse-variance weighted mean to determine the central value for each category and the reduced chi-squared statistic () to assess internal consistency. We explicitly quantify the tension between the Distance Ladder (Category 1) and the various subsets of Sound-Horizon-independent measurements, concluding with a comparison against the combined ensemble of all One-Step methods. We also compare our results with our previous analysis [Perivolaropoulos2024] to assess the evolution of the tension landscape.
@@ III.1 Methodology
For a dataset of independent measurements , the weighted mean and its standard error are calculated as:
@@ III.2 Weighted Mean Statistics by Category
The statistical results for the four categories defined in Section II are summarized in Table 7. For comparison, we also include the results from our previous analysis [Perivolaropoulos2024], which employed a binary classification scheme (Distance Ladder vs. One-Step).
@@
| 1. Distance Ladder | 29 | 0.74 | — | |
| 2. Local One-Step (CDM) | 33 | 0.85 | ||
| 3. Pure Local One-Step | 15 | 0.78 | ||
| 4. CMB Sound Horizon Free | 6 | 0.36 | ||
| Combined (2+3+4) | 54 | 0.85 | ||
| Corr.-adjusted (2+3+4) | 54 | — | ||
| Comparison with Perivolaropoulos (2024) [Perivolaropoulos2024]: | ||||
@@ III.2.1 Category 1: Distance Ladder ()
The 29 measurements comprising the Distance Ladder category yield a weighted mean of km s-1 Mpc-1. As shown in Figure 2, this category forms a sharp, narrow peak that is clearly separated from the other distributions.
The reduced chi-squared indicates good internal consistency, with the scatter being slightly smaller than expected from the reported uncertainties. This value is notably higher than our 2024 analysis (), suggesting that the expanded dataset—which now includes measurements with greater methodological diversity (e.g., JAGB, Miras, pairwise velocities)—exhibits more realistic scatter. Nevertheless, may indicate that: (i) reported uncertainties are somewhat conservative, (ii) positive correlations exist between measurements sharing common calibration anchors, or (iii) there is a tendency for published results to cluster around previously established values.
The high precision of this combined estimate is driven by the recent JWST Cepheid measurements [Riess2025] and compatible results from other calibrators including TRGB [Jensen2025], Miras [Bhardwaj2025], and SBF [Blakeslee2021]. We note that the Lee et al. [Lee2024] JAGB measurement () represents a outlier within this category, as discussed in Section II.
III.2.2 Category 2: Local One-Step CDM ()
@@
III.2.3 Category 3: Pure Local One-Step ()
The 15 model-independent measurements (Megamasers, Cosmic Chronometers, EPM) yield an intermediate value of km s-1 Mpc-1. Crucially, the removal of the CDM shape assumption results in a central value shift of km s-1 Mpc-1 relative to Category 2. This shift is visible in Figure 2 as a broader distribution positioned between the CDM-dependent results and the Distance Ladder.
The reduced chi-squared is lower than unity, indicating either conservative error estimates or the presence of correlations. This category exhibits significant internal heterogeneity: the Vogl et al. [Vogl2024] tailored EPM measurement () strongly favors the Distance Ladder value, while most Cosmic Chronometer measurements favor values closer to Planck. This heterogeneity suggests that Category 3 may itself contain distinct subpopulations, a possibility we explore further in Section IV.
III.2.4 Category 4: CMB Sound Horizon Free ()
The 6 measurements utilizing CMB data without the sound horizon yield km s-1 Mpc-1. This confirms that early-universe constraints on matter density (via the equality scale ) favor a low regardless of the calibration [Pogosian2024].
The very low reduced chi-squared warrants careful interpretation. With only measurements, statistical fluctuations in are expected to be large (the standard deviation of scales as for ). Nevertheless, this low value likely reflects: (i) the dominance of the two high-precision Pogosian et al. [Pogosian2024] measurements, which share a common methodology and partially overlapping datasets; and (ii) the fact that all measurements in this category are ultimately constrained by similar early-universe physics. In Figure 2, this distribution overlaps significantly with Category 2, reinforcing the consistency between early-universe physics and local data when CDM is assumed.
III.3 Comparison: Distance Ladder vs. Combined One-Step
To test the robustness of the “Early vs. Late” narrative, we combine all 54 Sound-Horizon-independent measurements from Categories 2, 3, and 4 into a single “All One-Step” ensemble. This combined sample represents the entirety of the evidence against the Distance Ladder that does not rely on the standard ruler.
@@ III.2.4 Category 4: CMB Sound Horizon Free ()
This distinction explains why the curves overlap visually even when the weighted means differ by several standard deviations. The tension between the Distance Ladder and the Combined One-Step sample is:
@@ III.3.1 Comparison with Previous Results
Table 7 includes a comparison with our 2024 analysis [Perivolaropoulos2024]. Despite the significant expansion of the dataset (from 53 to 83 measurements) and the refinement of the classification scheme (from 2 to 4 categories), the core results remain consistent:
@@
The One-Step weighted mean has shifted from to km s-1 Mpc-1, remaining consistent within uncertainties.
@@ III.3.2 Impact of Correlations
The tension quoted above assumes statistical independence between measurements. As detailed in Appendix C, accounting for realistic correlations within each category increases the combined uncertainties. Using correlation coefficients estimated from shared calibration data and common methodology (– for Distance Ladder; – for One-Step), we obtain:
@@ III.4 Internal Tension: Model Dependence
Beyond the external tension with the Distance Ladder, we identify a notable internal tension among the One-Step measurements based on model assumptions. Comparing Category 2 (Local One-Step CDM) with Category 3 (Pure Local One-Step), we find:
@@
Sample size: Category 3 contains only 15 measurements, making the weighted mean sensitive to individual high-precision entries.
@@ III.5 Summary of Statistical Results
The combined One-Step measurements yield km s-1 Mpc-1, consistent with Planck.
@@
Local measurements assuming CDM (Category 2) yield systematically lower than model-independent local measurements (Category 3), with a internal tension.
@@ IV The Impact of Model Assumptions
The results of Section III reveal a striking feature of the Hubble tension landscape: the discrepancy is not merely between the Distance Ladder and the Early Universe, but also manifests as an internal tension within the domain of One-Step measurements. Specifically, the statistical compatibility of a measurement with the Planck value depends critically on whether the standard CDM expansion history is assumed during the inference process. In this section, we dissect this model dependence, examine the critical role of specific high-precision measurements, and explore the physical implications of the observed bimodality in the distribution.
@@ IV.1 Implicit Bias of the CDM Expansion History
The most significant finding of our four-category classification is the tension between local One-Step measurements that assume a CDM cosmology (Category 2: km s-1 Mpc-1) and those that do not (Category 3: km s-1 Mpc-1).
Measurements in Category 2, such as Gravitational Wave standard sirens [Abbott2021] and standard analyses of Strong Lensing time delays [Shajib2023], typically span the redshift range . To extract from the observed luminosity distances or time delays, these methods must integrate the Hubble parameter :
@@
Our analysis shows that this CDM shape constraint systematically prefers lower values ( km s-1 Mpc-1), aligning almost perfectly with the Planck CMB inference [Planck2018]. This demonstrates a remarkable self-consistency of the CDM model: if one assumes the CDM background evolution holds true at late times, local data (excluding the Distance Ladder) yields results consistent with early-universe physics, even without using the Sound Horizon.
@@
CDM is correct, and Category 3 has larger systematics: Model-independent methods may suffer from unrecognized systematic errors that bias high. The lack of a shape constraint allows more freedom for systematics to propagate into the final result.
@@
The true deviates from CDM: If the actual expansion history differs from the CDM prediction at (e.g., due to a late-time gravitational transition, dynamical dark energy, or modified gravity [Perivolaropoulos2024]), forcing a CDM fit would bias the inferred downward to compensate for the shape mismatch.
A related possibility is that the tension reflects not a global modification of , but rather a localized change in the physics governing distance calibrators. Refs [Desmond:2019ygn, Desmond:2020wep, Marra2021, Ruchika2025, Perivolaropoulos:2025gzo] investigated whether a local variation in fundamental physics—such as a change in the fine-structure constant or gravitational strength within the calibration volume—could reconcile the SH0ES measurement with other probes without requiring modifications to the global expansion history. The current data do not definitively distinguish between these scenarios.
@@
Among the Category 3 measurements, the Vogl et al. [Vogl2024] tailored Expanding Photosphere Method (EPM) result deserves particular attention. This measurement yields km s-1 Mpc-1 (with systematic uncertainties — by the authors own estimate comparable to the statistical errors — taken into account) is fully consistent with the Distance Ladder and in tension with the Planck value—yet it is completely independent of both the Distance Ladder calibration chain and cosmological model assumptions.
@@ IV.3 The Megamaser Perspective
A similar pattern emerges from Megamaser measurements. The Barua et al. [Barua2025] frequentist analysis of the Megamaser Cosmology Project (MCP) sample yields km s-1 Mpc-1, again consistent with the Distance Ladder.
@@
In our classification, we assigned the Barua et al [Barua2025] measurement to Category 2 because it relies primarily on Keplerian disk modeling without adopting CDM-based peculiar velocity priors but relies on the CDM prediction for the angular diameter distance. Indeed, when is left as a free parameter rather than fixed to the Planck value, the best-fit drops to km s-1 Mpc-1 with uncertainties approximately five times larger, highlighting the model dependence of the headline result. In contrast, earlier analyses that adopted peculiar velocity corrections from the SH0ES flow model from Pesce et al. [Pesce2020] would be more appropriately classified as having some distance ladder dependence.
@@
The earlier Megamaser measurements by Kuo et al. [Kuo2013] ( km s-1 Mpc-1) and Gao et al. [Gao2015] ( km s-1 Mpc-1) yielded lower values, though with larger uncertainties. The evolution toward higher in recent Megamaser analyses may reflect improved disk modeling, better peculiar velocity treatments, or statistical fluctuations in small samples.
IV.4 Bimodality in the One-Step Distribution

### v2

@@ II.5 Category 3: Pure Local (Cosmological-Model Independent)
| Kuo [95] | 2013 | 68.0 | 9.0 | Megamasers (MCP) |
| Gao [66] | 2015 | 66.0 | 6.0 | Megamasers |
| Zhang [192] | 2022 | 65.9 | 3.0 | CC + HII |
| Du [55] | 2023 | 71.5 | 3.8 | TD Lenses + GRBs |
| Sneppen [171] | 2023 | 67.0 | 3.6 | EPM + Kilonovae |
| Li [101] | 2024 | 66.3 | 3.7 | Lensing Bias Correction |
| Vogl [182] | 2024 | 74.9 | 2.7 | SNe II (Tailored EPM) |
| Song [172] | 2025 | 70.4 | 6.9 | GW + SGL Calibration |
| Favale [60] | 2025 | 68.8 | 3.0 | Cosmic Chronometers |
We highlight that this category exhibits significant internal heterogeneity. Notably, the Vogl [182] tailored EPM measurement ( km s-1 Mpc-1) strongly favor the Distance Ladder value, while most cosmic chronometer measurements favor lower values. This heterogeneity will be discussed further in Section IV.
II.6 Category 4: CMB Sound-Horizon-Free (CMB–SHF)
The final category includes 9 measurements that utilize early-Universe data (primarily CMB) but specifically avoid the Sound Horizon as a standard ruler. Instead, these methods rely on alternative physical scales that are calibrated independently of pre-recombination acoustic physics:
BAO without priors: Analyses that use BAO peak positions but treat as a free parameter, constraining through the shape of rather than absolute calibration [144].
@@ II.6 Category 4: CMB Sound-Horizon-Free (CMB–SHF)
| Reese [147] | 2003 | 61.0 | 18.0 | SZ Effect |
| Philcox [137] | 2022 | 64.8 | BAO scale,Matter–rad. eq. | |
| Colaco [36] | 2023 | 67.2 | 6.1 | SZ + X-ray |
| Pogosian [144] | 2024 | 68.05 | 0.94 | BAO-pre DESI (No ) |
| Pogosiana [145] | 2025 | 69.37 | 0.65 | BAO-DESI DR2 (No ) |
| Bahr-Kalus [13] | 2025 | 65.2 | 5.6 | DESI Turnover Scale |
| Garcia [70] | 2025 | 70.03 | 0.97 | BAO-DESI DR2 + (No ) |
| Zaborowski [191] | 2025 | 70.8 | BAO-DESI + (No ) | |
| Krolewski [93] | 2025 | 69.0 | 2.5 | DESI + CMB (No ) |
a
Private communication.
@@ II.7 Summary of Classification
| 1 | Distance Ladder | 30 | Yes | No |
| 2 | Local (CDM) | 33 | No | Yes |
| 3 | Pure Local | 16 | No | No |
| 4 | CMB–SHF | 9 | No | Yes∗ |
@@ III Statistical Analysis and Results
In this section, we present a detailed statistical analysis of the measurements classified in Section II. We utilize the inverse-variance weighted mean to determine the central value for each category and the reduced chi-squared statistic () to assess internal consistency. We explicitly quantify the tension between the Distance Ladder (Category 1) and the various subsets of Sound-Horizon-Free measurements, concluding with a comparison against the combined ensemble of all Distance Ladder Independent/Sound Horizon Free (DLI–SHF) methods. We also compare our results with our previous analysis [134] to assess the evolution of the tension landscape.
@@ III.1 Methodology
For a set of independent measurements , the weighted mean and its standard error are calculated as:
@@
In cases of asymmetric errors, the following Equation is adopted for computations involving the mean or a unified error value:
@@ III.2 Weighted Mean Statistics by Category
The statistical results for the four categories defined in Section II are summarized in Table 7. For comparison, we also include the results from our previous analysis [134], which employed a binary classification scheme (Distance Ladder vs. Distnance Ladder Independent/Sound Horizon Free).
@@
| 1. Distance Ladder | 30 | 0.72 | — | |
| 2. Local (CDM) | 33 | 0.80 | ||
| 3. Pure Local | 16 | 0.77 | ||
| 4. CMB–SHF | 9 | 0.89 | ||
| Combined (2+3+4) | 58 | 0.95 | ||
| Corr.-adjusted (2+3+4) | 58 | — | ||
| Comparison with Perivolaropoulos (2024) [134]: | ||||
@@ III.2.1 Category 1: Distance Ladder ()
The 30 measurements comprising the Distance Ladder category yield a weighted mean of km s-1 Mpc-1. As shown in Figure 2, this category forms a sharp, narrow peak that is clearly separated from the other distributions.
The reduced chi-squared indicates good internal consistency, with the scatter being slightly smaller than expected from the reported uncertainties. This value is notably higher than our 2024 analysis (), suggesting that the expanded data—which now includes measurements with greater methodological diversity (e.g., JAGB, Miras, pairwise velocities)—exhibits more realistic scatter. Nevertheless, may indicate that: (i) reported uncertainties are somewhat conservative, (ii) positive correlations exist between measurements sharing common calibration anchors, or (iii) there is a tendency for published results to cluster around previously established values.
The high precision of this combined estimate is driven by the recent JWST Cepheid measurements [149] and compatible results from other calibrators including TRGB [83], Miras [20], and SBF [22]. We note that the Lee et al. [96] JAGB measurement () represents a outlier within this category, as discussed in Section II.
III.2.2 Category 2: Local CDM ( km s-1 Mpc-1)
@@
III.2.3 Category 3: Pure Local ()
The 16 cosmological-model-independent measurements (Megamasers, Cosmic Chronometers, EPM) yield an intermediate value of km s-1 Mpc-1. Crucially, the removal of the CDM shape assumption results in a central value shift of km s-1 Mpc-1 relative to Category 2. This shift is visible in Figure 2 as a broader distribution positioned between the CDM-dependent results and the Distance Ladder.
The reduced chi-squared is lower than unity, indicating either conservative error estimates or the presence of correlations. This category exhibits significant internal heterogeneity: the Vogl et al. [182] tailored EPM measurement () strongly favors the Distance Ladder value, while most Cosmic Chronometer measurements favor values closer to Planck. This heterogeneity suggests that Category 3 may itself contain distinct subpopulations, a possibility we explore further in Section IV.
III.2.4 Category 4: CMB Sound Horizon Free (CMB–SHF)()
The 9 measurements utilizing CMB data without the Sound Horizon yield km s-1 Mpc-1. This confirms that early-universe constraints on matter density (via the equality scale ) favor a low regardless of the calibration [144].
The reduced chi-squared is very close to unity, indicating good internal consistency among the measurements in this category. However, given that several measurements share common methodologies and partially overlapping datasets—particularly the Pogosian et al. [144, 145] and García Escudero et al. [70] results, which all employ Sound-Horizon-Free BAO analyses with DESI data—we examine potential correlations in Appendix C. In Figure 2, this distribution overlaps significantly with CMB–SHF measurements, reinforcing the consistency between early-universe physics and local data when CDM is assumed.
III.3 Comparison: Distance Ladder vs. Combined Distance Ladder Independent/Sound Horizon Free (DLI–SHF)
To test the robustness of the “Early vs. Late” narrative, we combine all 58 DLI–SHF from Categories 2, 3, and 4 into a single ensemble. This combined sample represents the entirety of the evidence against the Distance Ladder that does not rely on the standard ruler.
@@ III.2.4 Category 4: CMB Sound Horizon Free (CMB–SHF)()
This distinction explains why the curves overlap visually even when the weighted means differ by several standard deviations. The tension between the Distance Ladder and the Combined DLI–SHF sample is:
@@ III.3.1 Comparison with Previous Results
Table 7 includes a comparison with our 2024 analysis [134]. Despite the significant expansion of the sample (from 53 to 88 measurements) and the refinement of the classification scheme (from 2 to 4 categories), the core results remain consistent:
@@
The DLI–SHF weighted mean has shifted from to km s-1 Mpc-1, remaining consistent within uncertainties.
@@ III.3.2 Impact of Correlations
The tension quoted above assumes statistical independence between measurements. As detailed in Appendix C, accounting for realistic correlations within each category increases the combined uncertainties. Using correlation coefficients estimated from shared calibration data and common methodology (– for Distance Ladder; – for DLI–SHF), we obtain:
@@ III.4 Internal Tension: Model Dependence
Beyond the external tension with the Distance Ladder, we identify a notable internal tension among the Distance Ladder Independent/Sound Horizon Free(DLI–SHF) measurements based on model assumptions. Comparing Category 2 (Local CDM) with Category 3 (Pure Local), we find:
@@
Sample size: Category 3 contains only 16 measurements, making the weighted mean sensitive to individual high-precision entries.
@@ III.5 Summary of Statistical Results
The combined DLI–SHF measurements yield km s-1 Mpc-1, slightly elevated compared to our previous results but closer to Planck.
@@
Local measurements assuming CDM (Category 2) yield systematically lower than cosmological-model-independent local measurements (Category 3), with a internal tension.
@@ IV The Impact of Model Assumptions
The results of Section III reveal a striking feature of the Hubble tension landscape: the discrepancy is not merely between the Distance Ladder and the Early Universe, but also manifests as an internal tension within the domain of Distance Ladder Independent/Sound Horizon Free (DLI–SHF) measurements. Specifically, the statistical compatibility of a measurement with the Planck value depends critically on whether the standard CDM expansion history is assumed during the inference process. In this section, we dissect this model dependence, examine the critical role of specific high-precision measurements, and explore the physical implications of the observed bimodality in the distribution.
@@ IV.1 Implicit Bias of the CDM Expansion History
The most significant finding of our four-category classification is the tension between local DLI–SHF measurements that assume a CDM cosmology (Category 2: km s-1 Mpc-1) and those that do not (Category 3: km s-1 Mpc-1).
Measurements in Category 2, such as Gravitational Wave standard sirens [102] and standard analyses of Strong Lensing time delays [163], typically span the redshift range . To extract from the observed luminosity distances or time delays, these methods must integrate the Hubble parameter :
@@
Our analysis shows that this CDM shape constraint systematically prefers lower values ( km s-1 Mpc-1), aligning almost perfectly with the Planck CMB inference [141]. This demonstrates a remarkable self-consistency of the CDM model: if one assumes the CDM background evolution holds true at late times, local data (excluding the Distance Ladder) yields results consistent with early-universe physics, even without using the Sound Horizon [16].
@@
CDM is correct, and Category 3 has larger systematics: Cosmological-model-independent methods may suffer from unrecognized systematic errors that bias high. The lack of a shape constraint allows more freedom for systematics to propagate into the final result.
@@
The true deviates from CDM: If the actual expansion history differs from the CDM prediction at (e.g., due to a late-time gravitational transition, dynamical dark energy, or modified gravity [134, 77, 142, 119, 123, 7]), forcing a CDM fit would bias the inferred downward to compensate for the shape mismatch.
A related possibility is that the tension reflects not a global modification of , but rather a localized change in the physics governing distance calibrators. Refs [44, 45, 112, 154, 128] investigated whether a local variation in fundamental physics—such as a change in the fine-structure constant or gravitational strength within the calibration volume—could reconcile the SH0ES measurement with other probes without requiring modifications to the global expansion history.The current data do not definitively distinguish between these scenarios. A detailed comparison of late-transition versus smooth H(z) deformation models has been presented in [8]. Earlier reanalyses of SH0ES data exploring additional degrees of freedom also found evidence for such model-dependent effects [130].
@@
Among the Category 3 measurements, the Vogl et al. [182] tailored Expanding Photosphere Method (EPM) result deserves particular attention. This measurement yields km s-1 Mpc-1 (with systematic uncertainties — by the authors’ own estimate comparable to the statistical errors — taken into account) is fully consistent with the Distance Ladder and in tension with the Planck value—yet it is completely independent of both the Distance Ladder calibration chain and cosmological model assumptions.
@@ IV.3 The Megamaser Perspective
A similar pattern emerges from Megamaser measurements. The Barua et al. [17] frequentist analysis of the Megamaser Cosmology Project (MCP) sample yields km s-1 Mpc-1, again consistent with the Distance Ladder.
@@
In our classification, we assigned the Barua et al [17] measurement to Category 2 because it relies primarily on Keplerian disk modeling without adopting CDM-based peculiar velocity priors but relies on the CDM prediction for the angular diameter distance. Indeed, when is left as a free parameter rather than fixed to the Planck value, the best-fit drops to km s-1 Mpc-1 with uncertainties approximately five times larger, highlighting the model dependence of the headline result. In contrast, earlier analyses that adopted peculiar velocity corrections from the SH0ES flow model from Pesce et al. [136] would be more appropriately classified as having some distance ladder dependence.
@@
The earlier Megamaser measurements by Kuo et al. [95] ( km s-1 Mpc-1) and Gao et al. [66] ( km s-1 Mpc-1) yielded lower values, though with larger uncertainties. The evolution toward higher in recent Megamaser analyses may reflect improved disk modeling, better peculiar velocity treatments, or statistical fluctuations in small samples.
IV.4 Bimodality in the Distance Ladder Independent/Sound Horizon Free (DLI–SHF) Distribution